OpenStudy (anonymous):

A $7,000 motorcycle depreciates at 10% each year. Which value would represent the base of the exponential function that models this situation?

OpenStudy (anonymous):

@study100

OpenStudy (anonymous):

you know the equation for exponential decay?

OpenStudy (anonymous):

y=a x b^x

OpenStudy (anonymous):

I use to be able to do this I just forgot.

OpenStudy (anonymous):

Yes, since we're talking about rates, use this equation y= a(1-r) ^x were a= starting amount r= rate (percent) in decimals x= time

OpenStudy (anonymous):

y= 7000(1-0.1)^x then you can simplify to get the equation

OpenStudy (anonymous):

Now would time be 365 due to the year?

OpenStudy (anonymous):

365 days in 1 day. They don't decrease by days, but years. Your time is n years. they didn't tell us the years though. Just asking for the equation.

OpenStudy (anonymous):

and the base value.

OpenStudy (anonymous):

the base of this exponential function is 1-r so the base is 1-0.1 = 0.9

OpenStudy (anonymous):

the whole equation is y= 7000(0.9)^x

OpenStudy (anonymous):

OHHH okay

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

\(\color{salmon}{\large\heartsuit}\large\text{You're Most Welcome }\color{salmon}\heartsuit\) \(~~~~~~~~~~~~~~~\color{green}{\large\ddot\smile}\color{blue}{\large\ddot\smile}\color{salmon}{\large\ddot\smile}\color{blue}{\large\ddot\smile}\color{salmon}{\large\ddot\smile}\) \(~~~~~~~~~~~~~~~~~~~~\color{salmon}{\large\bigstar}\color{salmon}{\large\bigstar}\color{salmon}{\large\bigstar} \)